OpenStudy (anonymous):

Solve the differential equation. xy' - (x + 1)y = 0

OpenStudy (anonymous):

x(dy/dx) - (x+1)y = 0 x(dy/dx) = (x+1)y dy/dx = [ (x+1)/x ] y 1/y (dy) = (x+1)/x dx ::: integrate can yu go from there?

OpenStudy (anonymous):

somehow i got (x-1)/x

OpenStudy (anonymous):

as a final answeR?

OpenStudy (anonymous):

noo instead of (x+1)/x dx

OpenStudy (anonymous):

my final answer was y = -1/2Ce^x(x)^-2

OpenStudy (anonymous):

i have a feeling that's not right

OpenStudy (anonymous):

hmm not what i got, wait i'm gonna type what i have

OpenStudy (anonymous):

ok now i'm up to ln(y) = x + ln(x) + C.. not sure what to do next

OpenStudy (anonymous):

\[xy' - (x + 1)y = 0\]\[x{dy \over dx} - (x+1)y = 0\]\[x{dy \over dx} = (x+1)y\]\[{dy \over dx} = {(x+1) \over x} (y)\]\[{1 \over y} dy = {(x+1) \over x} dx\] Integrate:\[\ln|y|= \ln|x|+x +C\]\[e^{lny} = e^{lnx+x+C}=e^{lnx}*e^{x}*e^{C}\]\[y = A*x*e^{x}\] \[y = Axe^{x}\] that's what i have

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

I think soo.. what is A?

OpenStudy (anonymous):

e^C = constant a constant which i called A, you can called it whatever..

OpenStudy (anonymous):

but did yu understand everything?

OpenStudy (anonymous):

ohh gotcha.. i understand it.. thank you.. can you help me with other problems too?

OpenStudy (anonymous):

good. i sure can, just post them and i'll try

OpenStudy (anonymous):

\[\int\limits_{1/2}^{2} \] [(e^1/x) / (x^2)] dx

OpenStudy (anonymous):

\[\int\limits_{2}^{2}{{e^{1} \over x} \over x^{2}} dx\] is that what yu tried to write

OpenStudy (anonymous):

\[{e^{1/x} \over x^{2}}\]

OpenStudy (anonymous):

that didn't come out on my computer - maybe bc i'm using a mac. but its the integration bounded from 1/2 to 2

OpenStudy (anonymous):

and the function that you just wrote

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

do a product rule between e^(1/x) and x^(-2)

OpenStudy (anonymous):

can you do that ?

OpenStudy (anonymous):

i'm not really sure.. the integration of e^(1/x) is just e^(1/x)?

OpenStudy (anonymous):

-e^(1/x)*x^-1??

OpenStudy (anonymous):

the final answer comes up to be e^2 - sqrt(e) i think

OpenStudy (anonymous):

how did you get that?